# Mesh-Informed Neural Operator : A Transformer Generative Approach

### v2

@@ 2 Preliminaries
Neural operators. Neural operators learn mappings between function spaces, extending deep learning beyond the fixed-dimensional vectors handled by classical networks. This function space perspective is especially important for scientific computing that is governed by partial differential equations (PDEs) (Kovachki et al., 2023; Li et al., 2021). A core theoretical property of neural operator is discretization convergence (agnosticism): as the mesh of input function is refined, the prediction approaches the unique continuous solution (Kovachki et al., 2023). The Fourier Neural Operator (FNO) (Li et al., 2021) achieves this property via spectral-domain convolutions, achieving quasi-linear complexity on regular grids. Its global mixing excels at capturing long-range dependencies but may miss fine-scale detail (required for generative tasks) unless many Fourier modes are retained (Liu-Schiaffini et al., 2024).
For handling functions on irregular grids, Graph Neural Operator (GNO) (Gilmer et al., 2017; Li et al., 2020) remains a powerful candidate. GNO shares the message-passing functionality of Graph Neural Network (GNN), but differs in that the search radius for neighbors is defined in physical coordinates and made consistent across resolutions to guarantee discretization convergence. This results in GNO having a local receptive field, however it is computationally inefficient for certain tasks (Li et al., 2023; Liu-Schiaffini et al., 2024) and often struggles to learn the high-frequency information required for many generative models. This difficulty arises from the standard message-passing mechanisms, which tend to act as low-pass filters (Li et al., 2018). More recent work has focused on improved scalability and performance aspects of Neural Operators on irregular grids and non-rectangular physical domains, e.g., Geometry-Informed Neural Operator (GINO), Universal Physical Transformer (UPT), Latent Neural Operator (LNO), and Transolver (Li et al., 2023; Alkin et al., 2025; Wang & Wang, 2024; Wu et al., 2024). However, it remains unclear whether these architectures will also be effective for functional generative modeling on irregular grids (see Appendix C).
@@
Grid-Based Architectures. The modern U-Net architecture used in diffusion models (Rombach et al., 2022) has become a de-facto standard for generative tasks on grid-like data. These advanced U-Nets are distinct from the variant used in some earlier neural operators (Rahman et al., 2023), the U-Net can effectively capture multi-scale information through their hierarchical structure, skip connections, and integration of attention and residual blocks. Nevertheless, they remain constrained to regular grids due to their reliance on standard convolutional neural network kernels. Another prominent line of work involves Transformer-based architectures. Standard Vision Transformers (ViTs) and Diffusion Transformers (DiTs) (Dosovitskiy et al., 2021; Peebles & Xie, 2023), first "patchify" the input into non-overlapping blocks and then apply full self-attention, incurring $O(N_{\text{patch}}^{2})$ time complexity. For images, $N_{\text{patch}}\!\ll\!H\times W$ (Height and Width), yet the patchify operation creates discontinuities at block borders that can lead to checkerboard artifacts (Fang et al., 2022). Moreover, both ViT and DiT presuppose a fixed rectangular grid so that patchification is well defined.
@@
MINO [Ours] local $+$ global low + high
@@ 3 Methods
Problem Formulation. Within the flow matching paradigm, MINO takes three primary inputs: a noisy function $\mathbf{f_{t}}$, the set of its observation positions $\mathbf{f_{pos}}$, and the corresponding time step $\mathbf{t}$, along with any optional conditioning variables. Specifically, the input function $\mathbf{f_{t}}$ is represented by its values at $N_{\text{in}}$ discrete locations, $\mathbf{f_{pos}}=\{p_{i}\}_{i=1}^{N_{\text{in}}}$. These locations are a discretization of a continuous domain $D\subset\mathbb{R}^{P_{\text{dim}}}$. Note that $N_{\text{in}}$ can be different per sample. The codomain (channel) of $\mathbf{f}_{t}$ has dimension $f_{\text{dim}}$, resulting in an input tensor of function values with shape $[f_{\text{dim}},N_{\text{in}}]$. The output is another function (velocity field $\mathbf{v}_{t}$) at these same locations. In the context of flow matching or diffusion, the output dimension matches the input, so the target output also has a shape of $[g_{\text{dim}},N_{\text{in}}]$, where $g_{\text{dim}}=f_{\text{dim}}$.
@@
Geometry Encoder. First, we apply sinusoidal embeddings (Vaswani et al., 2023; Peebles & Xie, 2023) to the time step $\mathbf{t}$ and the observation positions $\mathbf{f_{pos}}$, yielding $\mathbf{t_{emb}}$ and $\mathbf{p_{emb}}$ respectively, as shown in Eq. 2. The input function $\mathbf{f}_{t}$ is then concatenated with the embedded positions $\mathbf{p_{emb}}$ along the codomain before being passed to a GNO layer; this maps the function from its potentially irregular discretization $\{p_{i}\}_{i=1}^{N_{\text{in}}}$ to a latent representation on a predefined regular grid of query points $\{p^{\text{query}}_{j}\}_{j=1}^{N_{\text{node}}}$. The GNO mapping is based on neighbor-searching within a fixed radius $r$, where $N_{\text{node}}\ll N_{\text{in}}$. Following this GNO layer, a linear layer projects its output $\mathbf{f_{t}^{E}}$ into $\mathbf{h_{0}^{E}}$ of shape $[L_{\text{dim}},N_{\text{node}}]$.
@@
$\mathbf{f_{t}^{E}}=\text{GNO}(\mathbf{f_{t}}+\mathbf{p_{emb}}),~{}~{}\mathbf{h_{0}^{E}}=\text{Linear}(\mathbf{f_{t}^{E}})$ (3)
@@
The output of the geometry encoder has a fixed size of $[L_{\text{dim}},N_{\text{node}}]$, which is independent of the manner in which the input function $\mathbf{f_{t}}$ is discretized.
@@
Cross-Attention Decoder. The decoder employs a different cross-attention mechanism from that of the encoder to enable mapping the latent representation back into the output velocity field at the original observation locations. Specifically, $\mathbf{f_{t}}$ is first processed by an MLP and then concatenated with $\mathbf{p_{emb}}$ to form $\mathbf{f_{t}^{D}}$. We stack $M_{2}$ cross-attention blocks where $\mathbf{f_{t}^{D}}$ serves as the $Q$. The $KV$ pair is taken from the output of the (optional) processor $\mathbf{h^{P}}$ (or $\mathbf{h_{M_{1}}^{E}}$ if the processor is omitted). The time embedding $\mathbf{t_{emb}}$ is used for conditioning. Finally, a LayerNorm and a linear transformation are applied to the output of the last attention block to produce the estimated velocity field $\mathbf{v_{t}}$.
@@
$\mathbf{h_{l}^{D}}=\text{MHCA}(Q=\mathbf{f_{t}^{D}},KV=\mathbf{h^{D}_{l-1}},C=\mathbf{t_{emb}}),\qquad l=1,\cdots,M_{2}$ (7)
$\mathbf{v_{t}}=\text{Linear}(\text{LayerNorm}(\mathbf{h_{M_{2}}^{D}})$ (8)
@@ 4 Experiments
In this section we empirically evaluate MINO and other baselines on a suite of functional generative benchmarks under OFM (Shi et al., 2025) paradigm due to its concise formulation and SOTA performance among functional generative paradigms. For each task we parameterize the velocity field $v_{\theta}(f_{t},t)$ with different neural-operator backbones to study how the choice of architecture affects performance. The benchmarks cover both regular and irregular grids, providing a comprehensive assessment. To the best of our knowledge, this is the first systematic comparison of modern neural-operator architectures on function-generation problems. We evaluate two variants of our proposed architecture: (i) MINO-U: a Diffusion U-Net as the latent space processor, and (ii) MINO-T: a pure Transformer-based variant without the latent space processor, where the encoder receives additional cross-attention blocks.
@@
Datasets. We curated a benchmark of six challenging functional datasets to evaluate performance across different domain types: (i) Rectangular Domains and Regular Grids: We use Navier-Stokes (Li et al., 2021), Shallow Water equation (Takamoto et al., 2024), and Darcy Flow (Takamoto et al., 2024) datasets. (ii) Irregular Domains and Grids: We use the Cylinder Flow dataset (Han et al., 2022), a synthetic Mesh-GP dataset on irregular meshes (Zhao et al., 2022), and a real-world global climate dataset (Dupont et al., 2022). These datasets are summarized in Table 2. Detailed descriptions of the datasets and their associated preparation are provided in Appendix A.
Baselines. To ensure a thorough comparison, we selected several state-of-the-art neural operators tailored to different domains : (i) For irregular grids, we benchmark against leading architectures including the Universal Physical Transformer (UPT) (Alkin et al., 2025), Latent Neural Operator (LNO) (Wang & Wang, 2024), Transolver (Wu et al., 2024), and Geometry-Informed Neural Operator (GINO) (Li et al., 2023). (ii) For regular grids, we further include the Fourier Neural Operator (FNO) as a strong, established baseline. To ensure a fair assessment, the parameter counts (or computational budgets) of all baseline models were carefully matched to our variants. It is important to note that, with the exception of FNO, all evaluated models (including our MINO variants) are designed to be domain-agnostic and process regular and irregular grids identically. A detailed description of the the baseline implementation is provided in Appendix A.
@@
Results. A visualization of the performance of MINO is provided in Figure 3 and as shown in Table 4, on regular grid benchmarks, our MINO variants achieve SOTA performance: MINO-U achieves best performance on the Navier-Stokes and Shallow Water datasets, while MINO-T excels on the Darcy Flow benchmark. In contrast, several baseline neural operators (GINO, LNO, UPT) perform poorly. Further analysis in Appendix CD suggests this is because these models primarily learn low-frequency information and fail to capture the high-frequency components essential for these generative tasks. FNO and Transolver achieve comparable, albeit inferior, results to our models. Notably, beyond superior accuracy, our MINO variants are also significantly more efficient than Transolver. As detailed in Table LABEL:tab:comptuation_vs, they require substantially less computation to train, are more GPU memory-efficient, and achieve up to a 2.9x speedup during inference, demonstrating the efficiency of the MINO architecture.
For tasks on irregular grids, our models continue to demonstrate top performance. MINO-U achieves the best results on the Cylinder Flow and Global Climate benchmarks, while MINO-T is the top performer on the Mesh-GP benchmark. Notably, MINO-U outperforms Transolver in all cases with less or equal computation (see Appendix A). Last, We perform an ablation study, which confirm the effectiveness the each component of our architecture, as detailed in Appendix E.
@@
MINO-T (ours) $3.0\cdot 10^{-2}$ $3.0\cdot 10^{-2}$ $\mathbf{6.0\cdot 10^{-2}}$ $\mathbf{2.3\cdot 10^{-2}}$ $3.5\cdot 10^{-2}$ $3.7\cdot 10^{-2}$
MINO-U (ours) $\mathbf{2.2\cdot 10^{-2}}$ $\mathbf{1.7\cdot 10^{-2}}$ $8.8\cdot 10^{-2}$ $3.4\cdot 10^{-2}$ $\mathbf{2.5\cdot 10^{-2}}$ $\mathbf{2.4\cdot 10^{-2}}$
@@
  MINO-T (Ours)    $4.0\cdot 10^{-2}$ $3.1\cdot 10^{-2}$ $\mathbf{3.9\cdot 10^{1}}$ $3.6\cdot 10^{-4}$ $5.6\cdot 10^{-5}$
  MINO-U (Ours)    $\mathbf{3.0\cdot 10^{-2}}$ $\mathbf{2.2\cdot 10^{-2}}$ $5.0\cdot 10^{1}$ $\mathbf{1.7\cdot 10^{-4}}$ $\mathbf{4.4\cdot 10^{-5}}$
@@
  FNO    $1.0\cdot 10^{-2}$ $9.4\cdot 10^{-3}$ $\mathbf{7.9\cdot 10^{-3}}$ $2.8\cdot 10^{-6}$ $1.0\cdot 10^{-4}$
  MINO-T (Ours)    $8.3\cdot 10^{-3}$ $6.9\cdot 10^{-3}$ $4.4\cdot 10^{-2}$ $5.9\cdot 10^{-6}$ $\mathbf{1.6\cdot 10^{-5}}$
  MINO-U (Ours)    $\mathbf{6.3\cdot 10^{-3}}$ $\mathbf{3.3\cdot 10^{-3}}$ $2.9\cdot 10^{-2}$ $\mathbf{6.3\cdot 10^{-7}}$ $2.2\cdot 10^{-5}$
Darcy Flow  GINO    $4.0\cdot 10^{-1}$ $3.0\cdot 10^{-1}$ $5.5\cdot 10^{2}$ $5.4\cdot 10^{-3}$ $1.8\cdot 10^{-3}$
@@
  MINO-T (Ours)    $\mathbf{7.5\cdot 10^{-2}}$ $\mathbf{2.3\cdot 10^{-2}}$ $\mathbf{4.7\cdot 10^{2}}$ $\mathbf{2.3\cdot 10^{-2}}$ $\mathbf{7.2\cdot 10^{-5}}$
  MINO-U (Ours)    $7.9\cdot 10^{-2}$ $4.8\cdot 10^{-2}$ $5.3\cdot 10^{2}$ $3.8\cdot 10^{-2}$ $2.0\cdot 10^{-4}$
Table 5: Computational efficiency and performance comparison between MINO variants and Transolver during training and inference on Shallow Water benchmark.
@@
MINO-U (ours) 19.2 M 0.419 GB 320 s 2.3×2.3$\times$2.3 ×
@@
Transolver 0.017 0.018 1.49 s 1×1$\times$1 ×
MINO-T (ours) 0.0083 0.0069 0.51 s $\mathbf{2.9\times}$
MINO-U (ours) $\mathbf{0.0063}$ $\mathbf{0.0033}$ 0.57 s 2.6×2.6$\times$2.6 ×

### v3

@@ 2 Preliminaries
Neural operators. Neural operators learn mappings between function spaces, extending deep learning beyond the fixed‑dimensional vectors handled by classical networks. This function space perspective is especially important for scientific computing that is governed by partial differential equations (PDEs) (Kovachki et al., 2023; Li et al., 2021). A core theoretical property of neural operators is discretization convergence (agnosticism): as the mesh of input function is refined, the prediction approaches the unique continuous solution (Kovachki et al., 2023). The Fourier Neural Operator (FNO) (Li et al., 2021) achieves this property via spectral‑domain convolutions, achieving quasi‑linear complexity on regular grids. Its global mixing excels at capturing long‑range dependencies but may miss fine‑scale detail (required for generative tasks) unless many Fourier modes are retained (Liu-Schiaffini et al., 2024).
For handling functions on irregular grids, Graph Neural Operator (GNO) (Gilmer et al., 2017; Li et al., 2020) remains a powerful candidate. GNO shares the message-passing functionality of Graph Neural Network (GNN), but differs in that the search radius for neighbors is defined in physical coordinates and made consistent across resolutions to guarantee discretization convergence. This results in GNO having a local receptive field, however it is computationally inefficient for certain tasks (Li et al., 2023; Liu-Schiaffini et al., 2024) and often struggles to learn the high-frequency information required for many generative models. This difficulty arises from the standard message-passing mechanisms, which tend to act as low-pass filters (Li et al., 2018; Giovanni et al., 2023). More recent work has focused on improved scalability and performance aspects of Neural Operators on irregular grids and non-rectangular physical domains, e.g., Geometry‑Informed Neural Operator (GINO), Universal Physical Transformer (UPT), Latent Neural Operator (LNO), and Transolver (Li et al., 2023; Alkin et al., 2024; Wang & Wang, 2024; Wu et al., 2024). However, it remains unclear whether these architectures will also be effective for functional generative modeling on irregular grids (see Appendix C).
@@
Grid-Based Architectures. The modern U-Net architecture used in diffusion models (Rombach et al., 2022) has become a de-facto standard for generative tasks on grid-like data. These advanced U-Nets are distinct from the variant used in some earlier neural operators (Rahman et al., 2023), the U-Net can effectively capture multi-scale information through their hierarchical structure, skip connections, and integration of attention and residual blocks. Nevertheless, they remain constrained to regular grids due to their reliance on standard convolutional neural network kernels. Another prominent line of work involves Transformer-based architectures. Standard Vision Transformers (ViTs) and Diffusion Transformers (DiTs) (Dosovitskiy et al., 2021; Peebles & Xie, 2023), first "patchify" the input into non‑overlapping blocks and then apply full self‑attention, incurring $O(N_{\text{patch}}^{2})$ time complexity. For images, $N_{\text{patch}}\!\ll\!H\times W$ (Height and Width), yet the patchify operation creates discontinuities at block borders that can lead to checkerboard artifacts (Fang et al., 2023). Moreover, both ViT and DiT presuppose a fixed rectangular grid so that patchification is well defined.
@@
MINO [Ours] local ++ global low + high
@@ 3 Methods
Problem Formulation. Within the flow matching paradigm, MINO takes three primary inputs: a noisy function $\mathbf{f_{t}}$, the set of its observation positions $\mathbf{f_{pos}}$, and the corresponding time step $\mathbf{t}$, along with any optional conditioning variables. Specifically, the input function $\mathbf{f_{t}}$ is represented by its values at $N_{\text{in}}$ discrete locations, $\mathbf{f_{pos}}=\{p_{i}\}_{i=1}^{N_{\text{in}}}$. These locations are a discretization of a continuous domain $D\subset\mathbb{R}^{P_{\text{dim}}}$. Note that $N_{\text{in}}$ can be different per sample. The codomain (channel) of $\mathbf{f}_{t}$ has dimension $f_{\text{dim}}$, resulting in an input tensor of function values with shape $f_{\text{dim}},N_{\text{in}}]$. The output is another function (velocity field $\mathbf{v}_{t}$) at these same locations. In the context of flow matching or diffusion, the output dimension matches the input, so the target output also has a shape of $g_{\text{dim}},N_{\text{in}}]$, where $g_{\text{dim}}=f_{\text{dim}}$.
@@
Geometry Encoder. First, we apply sinusoidal embeddings (Vaswani et al., 2017; Peebles & Xie, 2023) to the time step $\mathbf{t}$ and the observation positions $\mathbf{f_{pos}}$, yielding $\mathbf{t_{emb}}$ and $\mathbf{p_{emb}}$ respectively, as shown in Eq. 2. The input function $\mathbf{f}_{t}$ is then concatenated with the embedded positions $\mathbf{p_{emb}}$ along the codomain before being passed to a GNO layer; this maps the function from its potentially irregular discretization $\{p_{i}\}_{i=1}^{N_{\text{in}}}$ to a latent representation on a predefined regular grid of query points $\{p^{\text{query}}_{j}\}_{j=1}^{N_{\text{node}}}$. The GNO mapping is based on neighbor-searching within a fixed radius $r$, where $N_{\text{node}}\ll N_{\text{in}}$. Following this GNO layer, a linear layer projects its output $\mathbf{f_{t}^{E}}$ into $\mathbf{h_{0}^{E}}$ of shape $L_{\text{dim}},N_{\text{node}}]$.
@@
$\mathbf{f_{t}^{E}}=\text{GNO}(\text{concat}(\text{MLP}(\mathbf{f_{t}}),\mathbf{p_{emb}})),~~\mathbf{h_{0}^{E}}=\text{Linear}(\mathbf{f_{t}^{E}})$ (3)
@@
The output of the geometry encoder has a fixed size of $L_{\text{dim}},N_{\text{node}}]$, which is independent of the manner in which the input function $\mathbf{f_{t}}$ is discretized.
@@
Cross-Attention Decoder. The decoder employs a different cross-attention mechanism from that of the encoder to enable mapping the latent representation back into the output velocity field at the original observation locations. Specifically, $\mathbf{f_{t}}$ is first processed by an MLP and then concatenated with $\mathbf{p_{emb}}$ to form $\mathbf{f_{t}^{D}}$. In a single cross-attention block, $\mathbf{f_{t}^{D}}$ serves as the query ($Q$), while the key-value ($KV$) pair is taken from the output of the optional processor $\mathbf{h^{P}}$ (or $\mathbf{h_{M_{1}}^{E}}$ if the processor is omitted). The time embedding $\mathbf{t_{emb}}$ provides conditioning. Finally, a LayerNorm and a linear transformation are applied to the output of this attention block to produce the estimated velocity field $\mathbf{v_{t}}$.
@@
$\mathbf{h_{1}^{D}}=\text{MHCA}(Q=\mathbf{f_{t}^{D}},KV=\mathbf{h^{P}},C=\mathbf{t_{emb}})$ (7)
$\mathbf{v_{t}}=\text{Linear}(\text{LayerNorm}(\mathbf{h_{1}^{D}})$ (8)
@@ 4 Experiments
In this section we empirically evaluate MINO and other baselines on a suite of functional generative benchmarks under the OFM (Shi et al., 2025) paradigm due to its concise formulation and SOTA performance among functional generative paradigms. For each task we parameterize the velocity field $v_{\theta}(f_{t},t)$ with different neural-operator backbones to study how the choice of architecture affects performance. The benchmarks cover both regular and irregular grids, providing a comprehensive assessment. To the best of our knowledge, this is the first systematic comparison of modern neural-operator architectures on function-generation problems. We evaluate two variants of our proposed architecture: (i) MINO-U: a Diffusion U-Net as the latent space processor, and (ii) MINO-T: a pure Transformer-based variant without the latent space processor, where the encoder receives additional cross-attention blocks.
@@
Datasets. We curated a benchmark of six challenging functional datasets to evaluate performance across different domain types: (i) Rectangular Domains and Regular Grids: We use Navier-Stokes (Li et al., 2021), Shallow Water equation (Takamoto et al., 2022), and Darcy Flow (Takamoto et al., 2022) datasets. (ii) Irregular Domains and Grids: We use the Cylinder Flow dataset (Han et al., 2022), a synthetic Mesh-GP dataset on irregular meshes (Zhao et al., 2022), and a real-world global climate dataset (Dupont et al., 2022). These datasets are summarized in Table 2. Detailed descriptions of the datasets and their associated preparation are provided in Appendix A.
Baselines. To ensure a thorough comparison, we selected several state-of-the-art neural operators tailored to different domains : (i) For irregular grids, we benchmark against leading architectures including the Universal Physical Transformer (UPT) (Alkin et al., 2024), Latent Neural Operator (LNO) (Wang & Wang, 2024), Transolver (Wu et al., 2024), and Geometry-Informed Neural Operator (GINO) (Li et al., 2023). (ii) For regular grids, we further include the Fourier Neural Operator (FNO) as a strong, established baseline. To ensure a fair assessment, the parameter counts (or computational budgets) of all baseline models were carefully matched to our variants. It is important to note that, with the exception of FNO, all evaluated models (including our MINO variants) are designed to be domain-agnostic and process regular and irregular grids identically. A detailed description of the the baseline implementation is provided in Appendix A.
@@
Results. A visualization of the performance of MINO is provided in Figure 3 and as shown in Table 4, on regular grid benchmarks, our MINO variants achieve SOTA performance: MINO-U achieves best performance on the Navier-Stokes and Shallow Water datasets, while MINO-T excels on the Darcy Flow benchmark. In contrast, several baseline neural operators (GINO, LNO, UPT) perform poorly. Further analysis in Appendix CD suggests this is because these models primarily learn low-frequency information and fail to capture the high-frequency components essential for these generative tasks. FNO and Transolver achieve comparable, albeit inferior, results to our models. Notably, beyond superior accuracy, our MINO variants are also significantly more efficient than Transolver. As detailed in Table LABEL:tab:comptuation_vs, they require substantially less computation to train, are more GPU memory-efficient, and achieve up to a 3.0x speedup during inference, demonstrating the efficiency of the MINO architecture.
For tasks on irregular grids, our models continue to demonstrate top performance. MINO-U achieves the best results on the Cylinder Flow benchmark, while MINO-T is the top performer on the Mesh-GP and Global Climate benchmarks. Then, We perform an ablation study, which confirm the effectiveness the each component of our architecture, as detailed in Appendix E. Last, we provide additional experiments in Appendix F and G to further demonstrate the superior performance of MINO.
@@
MINO-T (ours) $2.9\cdot 10^{-2}$ $2.6\cdot 10^{-2}$ $\mathbf{4.1\cdot 10^{-2}}$ $\mathbf{3.0\cdot 10^{-3}}$ $\mathbf{2.1\cdot 10^{-2}}$ $\mathbf{1.4\cdot 10^{-2}}$
MINO-U (ours) $\mathbf{2.3\cdot 10^{-2}}$ $\mathbf{2.1\cdot 10^{-2}}$ $7.2\cdot 10^{-2}$ $4.1\cdot 10^{-2}$ $2.8\cdot 10^{-2}$ $2.6\cdot 10^{-2}$
@@
  MINO-T (Ours)    $4.0\cdot 10^{-2}$ $3.6\cdot 10^{-2}$ $1.9\cdot 10^{1}$ $\mathbf{6.8\cdot 10^{-5}}$ $1.6\cdot 10^{-5}$
  MINO-U (Ours)    $\mathbf{2.8\cdot 10^{-2}}$ $\mathbf{1.9\cdot 10^{-2}}$ $\mathbf{1.3\cdot 10^{1}}$ $9.2\cdot 10^{-5}$ $\mathbf{1.4\cdot 10^{-5}}$
@@
  FNO    $1.0\cdot 10^{-2}$ $9.4\cdot 10^{-3}$ $7.9\cdot 10^{-3}$ $2.8\cdot 10^{-6}$ $1.0\cdot 10^{-4}$
  MINO-T (Ours)    $9.8\cdot 10^{-3}$ $8.7\cdot 10^{-3}$ $1.6\cdot 10^{-2}$ $5.2\cdot 10^{-6}$ $2.0\cdot 10^{-5}$
  MINO-U (Ours)    $\mathbf{4.0\cdot 10^{-3}}$ $\mathbf{1.1\cdot 10^{-3}}$ $\mathbf{4.3\cdot 10^{-3}}$ $\mathbf{5.1\cdot 10^{-7}}$ $\mathbf{1.7\cdot 10^{-5}}$
Darcy Flow  GINO    $4.0\cdot 10^{-1}$ $3.0\cdot 10^{-1}$ $5.5\cdot 10^{3}$ $5.4\cdot 10^{-2}$ $1.8\cdot 10^{-3}$
@@
  MINO-T (Ours)    $\mathbf{8.9\cdot 10^{-2}}$ $\mathbf{3.4\cdot 10^{-2}}$ $\mathbf{9.1\cdot 10^{2}}$ $\mathbf{4.4\cdot 10^{-2}}$ $9.4\cdot 10^{-5}$
  MINO-U (Ours)    $9.5\cdot 10^{-2}$ $4.8\cdot 10^{-2}$ $1.0\cdot 10^{3}$ $7.7\cdot 10^{-2}$ $\mathbf{3.8\cdot 10^{-5}}$
Table 5: Computational efficiency and performance comparison between MINO variants and Transolver during training and inference on Navier-Stokes benchmark.
@@
MINO-U (ours) 19.2 M 0.419 GB 279 s $2.6\times$
@@
Transolver 0.053 0.051 1.49 s $1\times$
MINO-T (ours) 0.040 0.036 0.51 s $2.9\times$
MINO-U (ours) $\mathbf{0.028}$ $\mathbf{0.019}$ 0.49 s $\mathbf{3.0\times}$